# Cross-Cultural Expert-Level Art Critique Evaluation with Vision-Language Models

### v2

@@ 4 Dataset
To validate our framework, we use the Vulca-Bench corpus Yu et al. (2026). The complete corpus contains 6,804 matched pairs across 7 cultural traditions. However, we exclude the Mural tradition (109 pairs) from evaluation due to an insufficient number of expert annotations for reliable Tier III calibration, resulting on a 6-culture evaluation subset of 6,695 image-critique pairs covering a total of 165 culture-specific dimensions (see Appendix A for a detailed list of the dimensions). From this subset, we extract 294 stratified samples with bilingual expert-written critiques which we use as evaluation samples (Eval.) for our experiments. On top of that, a disjoint 450-sample human-scored (HScored) subset of art critiques (298 training / 152 held-out test) serves as the gold-standard $S_{h}$ for fitting and validating the isotonic calibration function $g$ in Tier III. See Table 2 for a complete overview of our data distribution.
Each image-critique pair consists of (i) an artwork image, (ii) artist and title metadata, and (iii) an expert bilingual critique (Chinese: $\sim$500 characters; English: $\sim$300 words) annotated with culture-specific dimensions. Critiques explicitly identify L1–L5 layer coverage through dimension-specific tags (e.g., CN_L1_D1 for Chinese Visual Colour, see Appendix A), enabling the automatic computation of cultural coverage metrics.
@@
The 450-sample of human-scored art critiques was annotated by 6 domain experts which assigned independent dual ratings to each art critique. Annotators scored based on our five Tier II dimensions (i.e., Coverage, Alignment, Depth, Accuracy, Quality) on a 1-5 scale. Inter-annotator agreement ($\kappa_{w}=0.39$) was fair Yu et al. (2025b), however, lower agreement on L3-L5 dimensions ($\kappa=0.35$) compared to L1-L2 ($\kappa=0.52$) evidenced the interpretive nature of this task, especially in deeper cultural dimensions such as philosophical aesthetics.
@@ 5.1 Experimental Setup
We evaluate 15 VLMs on art critique generation considering our evaluation subset consisting of 294 samples. These models cover five providers including both closed and open weight models: GPT-5, GPT-5.2, GPT-5-mini, GPT-4o, GPT-4o-mini, Claude-Sonnet-4.5, Claude-Opus-4.5, Claude-Haiku-3.5, Gemini-2.5-Pro, Gemini-3-Pro, Qwen3-VL-235B, Qwen3-VL-Flash, Llama4-Scout, GLM-4V-Flash, and DeepSeek-VL2.
We compress images to $\leq$3.75MB, generate bilingual L1-L5 critiques using a unified prompt, and compute Tier I metrics. Claude Opus 4.5 (Tier II judge) then returns five 1-5 dimension scores (Coverage, Alignment, Depth, Accuracy, Quality) in JSON format. Scores are then calibrated using a 450-sample human subset (298 train / 152 held-out). Our default evaluation metric is the calibrated Tier III score ($S_{\text{primary}}=S_{\text{II}}^{*}$), although Tier II dimension-specific scores are also reported for analytical purposes.
@@ 5.2 Judge Selection
To select our Tier II judge, we compared 8 different models. As shown in Table 3, OpenAI models provided higher scores (mean 4.1-4.5) while Claude models were stricter (3.4-4.0). Furthermore, we also explored a dual-judge configuration, observing unreliable results where judges exhibited systematic scale mismatches. More specifically, we observed strong ICC disagreement in cross-judge scenarios ranging from -0.5 (Claude-Opus-4.5 + GPT-5, $n=150$) to 0.12 (Claude-Sonnet + GPT-5, $n=150$), all below our 0.6 threshold. In the end, single-judge design with isotonic calibration yielded stable scores aligned with human judgements. Therefore, we selected Claude Opus 4.5 for three main reasons: (1) stable rank discrimination, (2) consistent cultural gap direction, and (3) absence of self-favouritism (i.e., it does not favor its own outputs).
@@ 5.3 Calibration Validation
Dimension MAE MAE (cal)
Coverage 0.391 0.412
Alignment 0.511 0.482
Depth 0.537 0.486
Accuracy 0.534 0.552
Quality 0.461 0.431
Avg. 0.457 0.433
$\Delta$ -5.2%
Table 4: Human calibration results (n=152).
Table 4 shows the human calibration (Tier III) results. Isotonic calibration is done for each of the dimensions independently and then averaged as a final score. Although some dimensions (i.e., Coverage and Accuracy) do not improve with calibration, we observed an overall improvement of 5.1% in the final aggregated score. Our calibrated evaluation framework then produces this final aggregated score as its output ($S_{\text{II}}^{*}$), without focusing on specific dimensions.
@@ 6 Results and Analysis
6.1 Overall VLM Performance (RQ1)
Model $S_{\text{II}}^{*}$ Cov Align Depth Acc Qual TI
@@ 6.1 Overall VLM Performance (RQ1)
Gemini-3-Pro 3.78 4.50 4.44 4.51 4.48 4.49 3.37
Llama4-Scout 3.70 4.50 4.36 4.48 4.37 4.46 3.83
GPT-5 3.65 4.44 4.39 4.42 4.40 4.43 3.18
Claude-Opus-4.5 3.65 4.50 4.44 4.50 4.46 4.49 3.24
GPT-5.2 3.64 4.50 4.40 4.51 4.41 4.50 3.33
Qwen3-VL-Flash 3.53 4.50 4.54 4.48 4.56 4.48 3.51
GPT-5-mini 3.50 4.50 4.38 4.50 4.43 4.49 3.21
@@
Std ($\sigma$) 0.03 0.13 0.10 0.13 0.08
Table 5: VLM $S_{\text{II}}^{*}$ performance with five Tier II dimensions (Cov=Coverage, Align=Alignment, Acc=Accuracy, Qual=Quality) and Tier I (TI) scores. The best results are in bold, the worst are underlined.
First, we analyse the overall VLM performance on our proposed evaluation framework. Table 5 shows a clear difference in performance when considering deeper cultural features among the 15 VLMs. Gemini-2.5-Pro (4.02) and Qwen3-VL-235B (3.99) are some of the best performing models, producing the most complete and deep critiques. On the other hand, Claude models remain competitive but do not achieve outstanding results (i.e., Sonnet 3.87, Opus 3.65). We can observe an overall spread of 0.78 points, revealing reasonable differences across different models. When looking at Tier II specific dimensions, we can see that Cultural Alignment and Factual Accuracy ($\sigma=0.13$) are the most discriminative dimension across models. For instance, we observe that DeepSeek-VL2 achieves reasonable Coverage but underperforms on Alignment/Accuracy, indicating weaker cultural grounding rather than surface-level fluency deficits.
We also observe that Tier II scores cluster in the 4.0-4.5 range ($\sigma=0.03$-$0.13$). This clustering reflects that (i) LLM judges have preference for polished language shaped by web-scale image-text pre-training, and (ii) ceiling effects in Coverage ($\sigma=0.03$), where most VLMs achieve >>80% dimension coverage and receive near-maximum scores. The discriminative power of our proposed framework lies in the higher-variance dimensions (i.e., Alignment and Accuracy, $\sigma=0.13$), which capture relevant L3-L5 cultural understanding features where VLMs show greater differences.
6.2 Tier I vs Tier II Gap (RQ2)
DCRCSACDSLQS1-10.5-0.500.50.5110.850.851.07-1.070.89-0.890.47-0.47Score Gap (TI - TII)
(a) Tier I-II gap. Positive = overestimate.
WE3.53.53.63.63.73.73.83.83.793.793.633.633.613.613.593.593.563.563.543.54Mean Score
(b) Cultural bias. Western highest.
Figure 3: (a) Tier I-II gap by metric. (b) Average performance scores by culture.
@@ 6.2 Tier I vs Tier II Gap (RQ2)
Table 6 compares Tier I metrics with the human-scored art critiques (n=450). All metrics show weak-to-moderate correlation ($r\in[0.27,0.53]$), confirming that they are not the best choice as individual evaluation metrics, but that can be useful as complementary signals. The information stemming from Tier I metrics can indicate critical limitations (e.g., superficial dimension coverage, lack of cultural depth) in an art critique, complementary to our calibrated judge output.
@@
6.3 Cultural Bias (RQ3)
Figure 3 (b) shows that in general, Western art critiques were given higher scores by the LLM judge in our Tier II evaluation. More specifically, we observed a gap between scores attributed to Chinese and Western art of -0.17 (Cohen’s d=0.65d=-0.65, $p<0.001$).
When looking specifically at our models, we observe that eleven out of the fifteen favor Western art critiques, with GPT-4o showing strongest bias ($\Delta=-0.52$) while Claude-Opus-4.5 being the most neutral ($\Delta=0.00$). We interpret this as an observed trend under our sampling and evaluation template, rather than a definitive claim about intrinsic model biases.
@@ 7 Discussion
Table 7 illustrates the advantages of the proposed Tri-Tier framework for art critique evaluation. A high linguistic quality and richness can mask limited cultural understanding. The framework provides additional discriminant features that were missing on previous dual-judge setups by combining automated metrics (Tier I), five-dimensional template-based single-judge evaluation (Tier II), and evaluation calibration with human expert references (Tier III). Specifically, Tier I indicators flag fluent-but-shallow outputs where high Quality masks low cultural Alignment, as illustrated by the failure cases in Table 7.
Our findings have direct implications on the development of VLMs for cultural tasks. First, VLMs consistently perform well at visual-technical levels (L1–L2) but show significant variance at cultural-philosophical levels (L3–L5), with coverage drops up to 30%. This gap suggests that explicit hierarchical cultural information should be incorporated during pre-training, allowing models to understand deeper cultural aspects. Our error analysis shows that a majority of low-scored critiques mainly failed on Alignment and Accuracy dimensions, confirming a lack of cultural understanding as their main limitation. Furthermore, the low Tier I-II agreement (ICC <0.5<0.5) highlights that automated metrics cannot be a substitute. The case study in Table 7 further supports that specialized cultural modules may be necessary to complement VLMs in culture-specific symbolism understanding and interpretation.
Table 7: VLM critique excerpts illustrating success vs. failure patterns. Success cases demonstrate L3-L5 cultural depth with technique terminology and historical context, while failure cases remain at surface-level description despite fluent language.
Image Artwork Q Critique Excerpt Analysis
[Uncaptioned image] Wang Jian
Landscape Album
96 “…rocks rendered with pima cun (hemp-fiber strokes), combined with short pima and jiesuo cun…embodying the Song-Yuan landscape tradition Success
L2: technique terms
L4: historical lineage
[Uncaptioned image] Michelangelo
Creation of Adam
99 “…this ‘near-touch yet not quite touching’ moment becomes the dramatic climax, symbolizing life not yet fully awakened” Success
L3: symbolic reading
L5: theological depth
[Uncaptioned image] Shen Zhou
Farewell at Jingjiang
49 “This work by Ming dynasty Shen Zhou uses ink technique…composition is reasonable, brushwork is exquisite…overall artistic conception is profound” Failure
Template: generic praise
✗ No specific technique
[Uncaptioned image] E.A. Hornel
Springtime
67 “Color relationships function on multiple levels…careful balance of forms establishes visual equilibrium” Failure
L1 only: visual description
✗ No Glasgow School context
@@ 8 Conclusion
We introduced a Tri-Tier framework for cross-cultural art critique evaluation built on an L1-L5 cultural-depth schema. Tier I provides four lightweight, reference-free automated metrics for lower-bound assessment, Tier II applies template-based scoring with a single primary judge across five dimensions (Coverage, Alignment, Depth, Accuracy, Quality), and Tier III calibrates the aggregated Tier II scores to human expert ratings via isotonic regression.
We evaluated 15 VLMs across six different cultural traditions on a 294-sample evaluation set, and calibrated with a 450-sample subset containing human-scored art critiques, observing significant variations of performance in cultural understanding across five different dimensions. The dimensions with major variability (i.e., Alignment and Accuracy) highlight the limitations of VLMs as art critique generators, indicating that cultural grounding and factual accuracy still represent major challenges in culture-related tasks. Furthermore, we also observed a biased tendency between Western and Chinese art critiques, assigning higher scores to Western samples and being consistent across judge models. Finally, we emphasize that the proposed evaluation framework is designed to assist, not replace, expert interpretation in culturally sensitive settings.
Our study, therefore, provides relevant results pointing to some of the main limitations of current VLMs in cultural tasks. Future work will focus on implementing these improvements, seeking to deepen cultural understanding and reduce biases between different cultures.

### v3

@@ 4 Dataset
To validate our framework, we use the Vulca-Bench corpus Yu et al. (2026), [^1] Dataset: https://huggingface.co/datasets/harryHURRY/vulca-bench our companion dataset paper that details the full annotation protocol and quality assurance process. The complete corpus contains 6,804 matched pairs across 7 cultural traditions. However, we exclude the Mural tradition (109 pairs) from evaluation due to an insufficient number of expert annotations for reliable Tier III calibration, resulting in a 6-culture evaluation subset of 6,695 image-critique pairs covering a total of 165 culture-specific dimensions (see Appendix D for a detailed list of the dimensions). From this subset, we extract 294 stratified samples with bilingual expert-written critiques which we use as evaluation samples (Eval.) for our experiments. On top of that, a disjoint 450-sample human-scored (HScored) subset of art critiques (295 training / 155 held-out test) serves as the gold-standard $S_{\text{h}}$ for fitting and validating the sigmoid calibration function $g^{*}$ in Tier III. The pool size was determined by available annotator capacity; the 65/35 train-test split balances calibration stability against held-out evaluation power (see Appendix A.3 for a sample-size sensitivity analysis). See Table 3 for a complete overview of our data distribution.
Each image-critique pair consists of (i) an artwork image, (ii) artist and title metadata, and (iii) an expert bilingual critique (Chinese: $\sim$500 characters; English: $\sim$300 words) annotated with culture-specific dimensions. Critiques explicitly identify L1–L5 layer coverage through dimension-specific tags (e.g., CN_L1_D1 for Chinese Visual Colour, see Appendix D), enabling the automatic computation of cultural coverage metrics.
@@
The 450-sample human-scored subset was annotated by domain experts via a balanced incomplete block design: three annotator pairs each scored ${\sim}100$ art critiques independently on the five Tier II dimensions (1–5 scale, 0.5-step increments), yielding 299 dual-rated items for agreement analysis. Overall inter-annotator agreement was moderate ($\kappa_{w}=0.43$, 95% CI $[0.39,0.45]$; Landis and Koch, 1977). Surface-oriented dimensions (Coverage, Quality) showed higher agreement ($\kappa_{w}=0.54$ $[0.48,0.58]$) than L3–L5 dimensions requiring cultural interpretation (Alignment, Depth, Accuracy: $\kappa_{w}=0.35$ $[0.31,0.39]$), evidencing the inherent subjectivity of deeper cultural understanding assessment. [^2] These $\kappa_{w}$ values measure agreement on ordinal rubric ratings (1–5 scale on Tier II dimensions), a harder task than the binary dimension-presence agreement ($\kappa=0.77$) reported in Yu et al. (2026). The two figures are not directly comparable.
@@ 5.1 Experimental Setup
We evaluate 15 VLMs on art critique generation using 294 evaluation samples spanning six cultural traditions, yielding 4,405 model–sample evaluations.33315 models $\times$ 294 samples = 4,410. Five instances (0.11%) were excluded: 4 malformed JSON responses (GLM-4V-Flash, 1.4% of its evaluations) and 1 empty critique (GPT-5, 0.3%). All remaining 13 models had zero failures. Imputing the excluded instances with a default score of 3.0 instead of excluding them does not change any model ranking. The models cover five providers (both closed and open weight): GPT-5, GPT-5.2, GPT-5-mini, GPT-4o, GPT-4o-mini, Claude-Sonnet-4.5, Claude-Opus-4.5, Claude-Haiku-3.5, Gemini-2.5-Pro, Gemini-3-Pro, Qwen3-VL-235B, Qwen3-VL-Flash, Llama4-Scout, GLM-4V-Flash, and DeepSeek-VL2.
Each model receives a compressed artwork image ($\leq$3.75MB) and generates a bilingual L1–L5 critique using a unified prompt. Tier I automated metrics are computed on the generated critique. Claude Opus 4.5 (Tier II judge) then scores the five rubric dimensions in reference-guided mode (§3.1). The aggregate Tier II score is calibrated via sigmoid calibration on the human-scored subset (§4). Our default evaluation metric is the calibrated score $S_{\text{II}}^{*}$; dimension-specific Tier II scores are reported for diagnostic purposes.
@@ 5.2 Judge Selection
To select our Tier II judge, we compared 8 different models. As shown in Table 4, OpenAI models provided higher scores (mean 4.1-4.5) while Claude models were stricter (3.4-4.0). Furthermore, we also explored a dual-judge configuration, observing unreliable results where judges exhibited systematic scale mismatches. More specifically, we observed strong ICC(2,1) disagreement in cross-judge scenarios ranging from 0.50-0.50 (Claude-Opus-4.5 + GPT-5, $n=150$) to $0.12$ (Claude-Sonnet + GPT-5, $n=150$), all below our 0.6 threshold. In the end, single-judge design with sigmoid calibration yielded stable scores aligned with human judgements. Therefore, we selected Claude Opus 4.5 for three main reasons: (1) stable rank discrimination, (2) consistent cultural gap direction, and (3) absence of self-favouritism (i.e., it does not favour its own outputs).
@@ 5.3 Calibration Validation
Score MAE MAE (cal)
Aggregate $S_{\text{II}}$ 0.454 0.446
$\Delta$ -1.7%
Table 5: Human calibration results (n=155, held-out). Tier III calibrates only the aggregate Tier II score.
Table 5 reports held-out human calibration (Tier III). The sigmoid function (Eq. 6) is fitted on the training split and applied only to the aggregate Tier II score on held-out data, yielding a 1.7% MAE reduction. Dimension-specific scores remain uncalibrated (§3). A comparison with isotonic regression and a sample-size sensitivity analysis are reported in Appendix A.3.
@@ 6 Results and Analysis
6.1 Cross-Cultural VLM Diagnosis (RQ3)
Model $S_{\text{II}}^{*}$ Cov Align Depth Acc Qual
@@ 6.1 Cross-Cultural VLM Diagnosis (RQ3)
Claude-Opus-4.5 4.09 4.37 3.90 3.99 3.65 4.27
GPT-5 4.00 4.23 3.48 4.04 3.85 4.08
GPT-5.2 3.97 4.17 3.53 3.97 3.66 4.19
Gemini-3-Pro 3.95 4.12 3.60 3.83 3.74 4.11
Llama4-Scout 3.67 4.21 3.48 3.36 2.96 4.10
GPT-5-mini 3.64 3.87 2.98 3.48 3.67 3.97
Claude-Haiku-3.5 3.61 3.80 3.18 3.46 3.35 4.05
@@
Std ($\sigma$) 0.33 0.48 0.56 0.35 0.24
Table 6: VLM $S_{\text{II}}^{*}$ performance with five Tier II dimensions (Cov=Coverage, Align=Alignment, Depth=Critique Depth, Acc=Accuracy, Qual=Quality). The best results are in bold, the worst are underlined. $S_{\text{II}}^{*}$ is computed as $\sigma(\bar{S}_{\text{II}})$, i.e., sigmoid calibration applied to the per-model mean of raw dimension averages. Culture-level analyses (Figure 3b) instead average per-evaluation calibrated scores (difference <0.05<0.05).
Having established the instrument’s measurement properties (§5), we now apply it to diagnose VLM cultural understanding. Table 6 reveals clear performance stratification across cultural depth. Gemini-2.5-Pro (4.27) and Qwen3-VL-235B (4.21) produce the most complete critiques, while Claude models remain competitive (Sonnet 4.11, Opus 4.09). The overall spread of 1.26 points reveals meaningful differences across models. The most discriminative dimension is Critique Depth ($\sigma=0.56$), followed by Cultural Alignment ($\sigma=0.48$) and Factual Accuracy ($\sigma=0.35$), precisely the dimensions that correspond to iconographic and iconological interpretation (L3–L5). DeepSeek-VL2, for instance, scores 3.50 on Coverage but only 2.74 on Alignment and 2.64 on Depth, indicating weaker cultural grounding rather than surface-level fluency deficits. Bootstrap 95% confidence intervals (10K resamples) show that 6 of 14 adjacent-rank pairs are non-overlapping, while the top tier (#1–3: $S_{\text{II}}^{*}$ 4.11–4.27) and bottom tier (#13–15: 3.01–3.24) are fully separated ($p<0.001$, permutation test). Mid-table ranks with overlapping CIs should be interpreted as performance bands rather than strict orderings.
Linguistic Quality shows the least variance ($\sigma=0.24$), confirming that all 15 VLMs produce fluent text and that surface polish is not the bottleneck. The discriminative power lies in Depth and Alignment, which capture L3–L5 cultural understanding where pre-iconographic competence no longer suffices. This is consistent with the theoretical distinction in §2: mentioning relevant concepts is relatively easy, while correctly attributing and interpreting cultural properties is where models diverge.
6.2 Automated vs. Judge-Based Measurement (RQ2)
DCRCSACDSLQS1-10110+0.85-1.07-0.89-0.47  Overest.   Underest.Score Gap (TI - TII)(a) Tier I–II gap by metricWEINJPISKRCN3.63.63.83.8443.983.753.713.683.603.59Mean $S_{\text{II}}^{*}$(b) Mean $S_{\text{II}}^{*}$ by culture
Figure 3: (a) Tier I–II gap: positive = Tier I overestimates, negative = Tier I underestimates relative to Tier II. (b) Average calibrated scores by culture; Western art consistently receives highest scores.
@@
Table 7 compares Tier I metrics with human-scored art critiques ($n=450$). All metrics show weak-to-moderate correlation ($r\in[0.27,0.53]$), confirming their role as complementary risk indicators rather than standalone evaluation metrics.
The near-zero ICC for DCR (0.02) confirms that lexical coverage and semantic understanding operate as largely independent dimensions rather than coarse and fine-grained measures of the same construct.
@@
6.3 Cultural Disparities (RQ3)
The instrument also reveals culture-stratified scoring patterns. Figure 3 (b) shows that Western art critiques consistently receive higher scores in our Tier II evaluation. The gap between Chinese and Western art scores is 0.39-0.39 (Cohen’s d=0.74d=-0.74, computed over $n_{\text{CN}}{=}750$, $n_{\text{WE}}{=}747$ instance-level evaluations; $p<0.001$, bootstrap 95% CI [0.44,0.34][-0.44,-0.34], 1,000 resamples).
At the model level, thirteen of fifteen VLMs favour Western art critiques, with GPT-4o-mini showing strongest bias ($\Delta=-1.08$) and GPT-5.2 most neutral ($\Delta=+0.07$). Appendix Figure 4 visualizes the model-stratified gap distribution.
Confound controls.
Two analyses rule out major confounds. A genre-controlled landscape subset ($n_{\text{CN}}{=}300$, $n_{\text{WE}}{=}405$) yields a larger gap (d=0.93d{=}{-}0.93, 95% CI [1.09,0.78][-1.09,-0.78], $p<0.001$), exceeding the overall effect. A blind-culture pilot (n=50n{=}50, GPT-4o) shows the gap widens when the culture tag is removed ($\Delta_{\text{blind}}{=}{-}0.61$ vs. $\Delta_{\text{std}}{=}{-}0.54$), ruling out judge bias (Appendix A.5).
Appendix A provides additional robustness validation, including a Tier I–Tier II ICC analysis confirming that automated metrics and judge scores measure distinct constructs (all ICC <0.2<0.2).
@@ 7 Discussion
As illustrated by representative case studies (Table 14, Appendix E), high linguistic quality can mask limited cultural understanding. The tri-tier framework addresses this by combining automated risk indicators (Tier I) with rubric-based judge evaluation (Tier II) and human calibration (Tier III), providing discriminant features missing in previous dual-judge setups.
Failure pattern analysis.
Case studies (Table 14, Appendix E) reveal that high keyword coverage can mask shallow understanding: models cite cultural terms generically (e.g., “ink technique,” “artistic conception”) without specifying how they manifest in the target painting. Successful critiques instead identify specific techniques and contextualise them within art-historical traditions. This pattern confirms that the Tier I–II gap reflects a structural difference in cultural understanding depth.
Implications for cultural AI.
The L1–L2 to L3–L5 degradation pattern (§6.1) suggests that VLMs trained on image-caption pairs acquire descriptive competence but lack cultural grounding for deeper interpretation. Furthermore, the low Tier I–II agreement (ICC <0.5<0.5) confirms that automated metrics cannot substitute for semantic evaluation. These findings carry practical consequences for museums, art education platforms, and cultural heritage projects, all of which risk systematic misinterpretation of non-Western traditions without validated evaluation instruments.
@@ 8 Conclusion
We return to the three research questions posed in §1.
RQ1 (Measurement).
Cultural understanding beyond perception can be measured through a tri-tier protocol combining automated indicators, single-judge rubric scoring across 165 culture-specific dimensions, and human calibration, grounded in the art-theoretical hierarchy defined in §2.
RQ2 (Reliability).
Automated proxies and LLM-judge scoring measure different constructs: Tier I–II ICC remains below 0.2 for all four metrics, and cross-judge ICC reaches 0.50-0.50. Sigmoid calibration to human ratings via a single judge provides a more reliable alternative.
RQ3 (Diagnosis).
Under the validated instrument, cultural understanding degrades from L1–L2 to L3–L5, with Depth ($\sigma{=}0.56$) and Alignment ($\sigma{=}0.48$) as the most discriminative dimensions. Thirteen of fifteen models assign higher scores to Western art (Cohen’s d=0.74d=-0.74, $p<0.001$), though cross-judge scale incompatibility precludes definitive multi-judge confirmation.
Broader impact.
[…]
Acknowledgements
We thank Ramon Ruiz-Dolz for initial discussions on project framing, early-stage conceptual feedback, and manuscript polishing during the formative phase of this work.